OpenStudy (anonymous):

Use elimination to find the solution to the system of equations. 4x + y = 12 2x - 3y = 6 A. x = 3 y = 0 B. x = 0 y = 12 C. x = 7 y = -16 D. x = 27/2 y = 7

OpenStudy (phi):

try multiplying the first equation by 3 (that means multiply every term on both sides of the = sign by 3) what do you get ?

OpenStudy (anonymous):

Here is what I have so far: 3(4x + y = 12) 12x +3y = 36 2x - 3x = 6 ------------ 14x = 42 --- --- 0 14 42/14 = 3 Y=3

OpenStudy (phi):

pretty close. once you get 14x = 42 you divide both sides by 14 to get \[ \frac{14x}{14} = \frac{42}{14} \] 14/14 on the left is 1 and 1*x is just x so you get x= 42/14 x= 3 so it is x=3 (not y) can you find y using 4x + y = 12 (and x=3) ?

OpenStudy (anonymous):

AH! Okay! Thank you!! That is where I messed up! 4x + y = 12 When x =3... 12x + y = 12 12/12 = 0 12/12 = 0 Y = 0

OpenStudy (phi):

Your algebra looks wrong. you start with 4x + y = 12 When x =3... x=3 means we replace any x we see in the equation with 3. Look for x, erase it, put in a 3 instead: 4*3 + y= 12 (4x means 4*x, and with x=3, 4x means 4*3) 4*3 is 12, so this is 12+y= 12 can you "solve for y" ? (I would add -12 to both sides)

OpenStudy (anonymous):

OH!! So... 12+y=12 And then add -12? 12 + (-12) = 0 Now Y=0 :)

OpenStudy (phi):

closer. you add -12 to both sides like this 12+y=12 12-12 + y= 12-12 (notice we still have y. it does not go away) now simplify: 12-12 is 0 so we get 0 + y= 0 and 0+y is y (adding zero to y does nothing) you get y= 0

OpenStudy (anonymous):

AH! Okay, I thought you meant actually add -12! Thanks so much for your help!!!

OpenStudy (phi):

I did mean add -12, but to both sides so you should show the work for adding -12 to both sides if you have time, watch this video and the few after to really get the knack of this http://www.khanacademy.org/math/algebra/solving-linear-equations-and-inequalities/equations_beginner/v/simple-equations

OpenStudy (anonymous):

Will do.. Thanks again. (: